OpenStudy (anonymous):

Simplify the raidcal. 7 square root 72

OpenStudy (anonymous):

\[7\sqrt{72}=7 \sqrt{2^3 3^2}=42 \sqrt{2} \]

OpenStudy (raden):

alternative : 7 sqrt(72) = 7 sqrt(36 * 2) = 7 sqrt(36) sqrt(2) = 7(6)sqrt(2) = 42 sqrt(2)